INTRACAVITY LASER ABSORPTION SPECTROSCOPY OF PLATINUM NITRIDE IN THE NEAR IN-FRARED

\underline{LEAH C. O’BRIEN}, KAITLIN A. WOMACK, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652; JAMES J. O’BRIEN, SEAN WHITTEMORE, Department of Chemistry, University of Missouri, St Louis, MO 63121-4499.

The (2,0) band of the A^{2}Σ^{−} - X^{2}Π_{1/2} electronic transition of PtN has been recorded using intracavity laser absorption spectroscopy. Transitions from ^{194}P tN, ^{195}P tN, and ^{196}P tN isotopologues are observed, as well as the nuclear hyperfine splitting due to ^{195}P t with I=1/2. The results of the analysis will be presented and compared with ab initio calculations.